OpenStudy (anonymous):

sum of the series 2^(1/n)

terenzreignz (terenzreignz):

To infinity?

OpenStudy (anonymous):

\[\sum_{1}^{\infty} \sqrt[n]{2}\]

terenzreignz (terenzreignz):

Well, does this sum even exist? Remember, for a series \[\large \sum_{n=1}^{\infty}a_n\]to be convergent, it is necessary for the underlying sequence to converge to zero... \[\huge \lim_{n \rightarrow \infty}a_n=0\]